# Learning to (Learn at Test Time): RNNs with Expressive Hidden States

### v2

@@ Abstract
Self-attention performs well in long context but has quadratic complexity. Existing RNN layers have linear complexity, but their performance in long context is limited by the expressive power of their hidden state. We propose a new class of sequence modeling layers with linear complexity and an expressive hidden state. The key idea is to make the hidden state a machine learning model itself, and the update rule a step of self-supervised learning. Since the hidden state is updated by training even on test sequences, our layers are called Test-Time Training (TTT) layers. We consider two instantiations: TTT-Linear and TTT-MLP, whose hidden state is a linear model and a two-layer MLP respectively. We evaluate our instantiations at the scale of 125M to 1.3B parameters, comparing with a strong Transformer and Mamba, a modern RNN. Both TTT-Linear and TTT-MLP match or exceed the baselines. Similar to Transformer, they can keep reducing perplexity by conditioning on more tokens, while Mamba cannot after 16k context. With preliminary systems optimization, TTT-Linear is already faster than Transformer at 8k context and matches Mamba in wall-clock time. TTT-MLP still faces challenges in memory I/O, but shows larger potential in long context, pointing to a promising direction for future research.
@@
Figure 2: Comparing to Mamba, TTT-Linear has better perplexity and fewer FLOPs (left), and better use of long context (right). Evaluations follow Kaplan et al. [40]. Left: Scaling trends on Books, zoomed in between 350M and 1.3B parameters. At 760M and 1.3B, TTT-Linear outperforms Mamba in perplexity using fewer FLOPs, and outperforms Transformer under linear interpolation. Right: Transformer and TTT-Linear can keep reducing perplexity as they condition on more tokens, while Mamba cannot after 16k context. All methods have matched training FLOPs as Mamba 1.4B. Details in Subsection 3.2.
Figure 3: Forward time per token (latency) for batch size 16 as context length varies. All models are 1.3B (1.4B for Mamba). Forward time per token grows linearly for Transformer as context length increases, but stays roughly constant for the other two methods. TTT-Linear is faster than Transformer at 8k context and matches Mamba. Details in Subsection 3.3.
@@ 1 Introduction
The difficulty with long context is inherent to the very nature of RNN layers: Unlike self-attention, RNN layers have to compress context into a hidden state of fixed size. As a compression heuristic, the update rule needs to discover the underlying structures and relationships among thousands or potentially millions of tokens. In this paper, we begin with the observation that self-supervised learning can compress a massive training set into the weights of a model such as an LLM, which often exhibits deep understanding about the semantic connections among its training data – exactly what we need from a compression heuristic.
@@
Motivated by this observation, we design a new class of sequence modeling layers where the hidden state is a model, and the update rule is a step of self-supervised learning. Because the process of updating the hidden state on a test sequence is equivalent to training a model at test time, this new class of layers is called Test-Time Training (TTT) layers. We introduce two simple instantiations within this class: TTT-Linear and TTT-MLP, where the hidden state is a linear model and a two-layer MLP, respectively. TTT layers can be integrated into any network architecture and optimized end-to-end, similar to RNNs layers and self-attention.
@@
While the TTT layer is already efficient in FLOPs, we propose two practical innovations to make it efficient in wall-clock time. First, similar to the standard practice of taking gradient steps on mini-batches of sequences during regular training for better parallelism, we use mini-batches of tokens during TTT. Second, we develop a dual form for operations inside each TTT mini-batch, to better take advantage of modern GPUs and TPUs. The dual form is equivalent in output to the naive implementation, but trains more than 5×5$\times$5 × faster. As shown in Figure 3, TTT-Linear is faster than Transformer at 8k context and matches Mamba.
Evaluations and open problems.
While we have highlighted some results for TTT-Linear at the beginning of the paper, Section 3 presents more comprehensive evaluations for both TTT-Linear and TTT-MLP, and open problems exposed by our evaluations. For example, our evaluations following the Chinchilla recipe [34] do not cleanly fit a linear scaling trend even for the Transformer baseline. Constrained by our academic resources, we encourage the community to join us in exploring solutions to these problems.
Summary of contributions.
  1. 1.
    We propose TTT layers, a new class of sequence modeling layers where the hidden state is a model, and the update rule is self-supervised learning. Our perspective that the forward pass of a layer contains a training loop itself opens up a new direction for future research.
  2. 2.
    TTT-Linear, one simple instantiation of TTT layers, outperforms Transformers and Mamba in our evaluations ranging from 125M to 1.3B parameters.
  3. 3.
    We improve the hardware efficiency of TTT layers through mini-batch TTT and the dual form, making TTT-Linear already a practical building block for LLMs.
@@ 3.2 Long context: Books
To evaluate capabilities in long context, we experiment with context lengths ranging from 1k to 32k in 2×2$\times$2 × increments, using a popular subset of the Pile called Books3. The training recipe here is the same as for the Pile, and all experiments for the TTT layers are performed in one training run.121212 Following the Mamba paper, we always use 0.5M tokens per training batch regardless of context length. That means for context length $T$ we have $0.5\text{M}\,/\,T$ sequences per batch (assume divisible). From the subset of results in Figure 12, we make a few observations:
@@
Figure 13: Experiments on Books with context lengths ranging from 1k to 32k. We treat context length as a hyper-parameter and connect the selected points. Since we have Transformers trained from scratch and finetuned, we label them as TF pretrain and TF finetune. The left panel of Figure 2 is a zoomed-in view between 350M and 1.3B parameters, where Transformer is TF finetune, the stronger Transformer baseline.
Context length as a hyper-parameter.
While the length of the input sequence is determined by the user, the context length at which the language model processes the input is determined by the engineer as a design choice. Therefore, context length is a hyper-parameter that can be selected just as other ones.131313 To be precise, there are two hyper-parameters: the context length at which the LLM is trained, and one at which the LLM is evaluated. Both of them can be different from the sequence length, which is determined by the user. Transformers tend to perform poorly when the evaluation context is longer than the training context [18]. Therefore, we always evaluate at the training context length, making the two hyper-parameters the same. For LLMs with linear complexity, we select the argmin in perplexity, since every context length has the same FLOPs. For Transformers, longer context costs more FLOPs, so we form a convex hull of all the points in the log-log plot and connect those on the boundary.
From Figure 13, we make a few observations:
  • The lines of TTT-Linear and TTT-MLP, the best-performing methods, almost completely overlap. The lines of Mamba and TF finetune also mostly overlap after $10^{20}$ FLOPs.
  • TF finetune performs significantly better than TF pretrain, as it benefits from long context without incurring extremely large cost in training FLOPs. Note that the inference FLOPs of TF finetune and pretrain are equally poor, which is not reflected in this plot.
  • For all methods trained from scratch (including TF pretrain), perplexity becomes worse once the context length becomes too large. This trend is highlighted in Figure 19 (in Appendix). We leave further investigation of this trend to future work.
The left panel of Figure 2 is a zoomed-in view of Figure 13. For clarity, we leave TF pretrain out of Figure 2 and only show TF finetune (labeled as Transformer) since it is the stronger baseline. Figure 14 reproduces the right panel of Figure 2, now with TTT-MLP and additional discussion.
Figure 14: Reproduction of the right panel of Figure 2, now with TTT-MLP. Comparing to TTT-Linear, TTT-MLP performs slightly worse at short context but better at long context. This observation matches our expectation that the MLP as hidden state is more expressive than the linear model. Again, all methods have matched training FLOPs as Mamba 1.4B. For TTT-Linear and TTT-MLP, this protocol implies matched inference FLOPs. Transformer (TF finetune) has 2.8×2.8$\times$2.8 × the inference FLOPs, giving it an advantage as our baseline.
@@ 3.3 Wall-clock time
Figure 15: Benchmark on an NVIDIA A100 GPU with 80G HBM and PCIe connections. Left: Same as Figure 3, but now with TTT-MLP. Note that our Transformer baseline is significantly faster then that in the Mamba paper, because we use vLLM [46], a state-of-the-art serving system, instead of the HuggingFace Transformer [77]. Right: For generate (decode), TTT-Linear and Mamba have almost the same latency, which is significantly smaller than that of Transformer and TTT-MLP.
@@
Due to resource constraints, our experiments are written in JAX and run on TPUs. On a v5e-256 TPU pod, the Transformer baseline takes 0.30s per iteration of training at context 2k, while TTT-Linear takes 0.27s per iteration, already 10% faster without any systems optimization. However, Mamba (implemented in PyTorch, Triton, and CUDA) can only run on GPUs, so for fair comparison, we also rewrite our method with preliminary systems optimization to run on GPUs.
Specifically, we write a GPU kernel for forward in ThunderKittens [66]. Historically, RNNs have been inefficient during forward and backward due to poor use of parallelism and matmuls. Our goal with the forward kernel is to demonstrate the effectiveness of mini-batch TTT and the dual form for these problems. A kernel for backward should have the same properties in efficiency as forward, but requires more complex logic for manual differentiation, therefore is left for future work.
The left panel of Figure 15 shows the latency for batch size 16 of our forward kernel. All models are 1.3B (1.4B for Mamba). Time per token grows linearly for Transformer as context length increases, but stays roughly constant for the other methods.141414 We observe that forward latency of the network increases slightly for TTT-Linear, TTT-MLP, and Mamba, even though latency of each sequence modeling layer alone stays constant. Consider the operation $\theta X$, where $\theta$ is $d\times d$ and $X$ is $d\times T$. Its latency (normalized over $T$) is expected to be constant, but in practice grows slightly with $T$. One possible cause of this phenomenon is the GPU throttling after $T$ gets very large [30]. Note that our Transformer baseline is significantly faster that in the Mamba paper, because we use vLLM [46], a state-of-the-art serving system, instead of the HuggingFace Transformer [77].
In addition, we write another GPU kernel for generate in Triton [72], and benchmark its speed in the right panel of Figure 15 for batch size 512. Another popular metric for wall-clock time is throughput, which takes into account the potential benefit of being able to use a larger batch size.151515 To calculate throughput for each method, we increase its batch size in 2×2$\times$2 × increments until GPU runs out of memory, measure the tokens per second for every batch size, and select the highest. For completeness, we report the throughput for forward and generate in Figure 20 (in Appendix). All the observations and ordering between methods above still hold for throughput.
@@ 4 Related Work
4.1 Modern RNNs
Mamba is one of the many Structured State-Space Models [27, 21, 57, 18]. The hidden state in these models is a vector, similar to in LSTMs. For TTT-Linear or TTT-MLP, the hidden state is a matrix or two matrices, therefore larger. In Figure 14, we find that TTT layers can take advantage of their larger hidden states to compress more information in long context, where TTT-MLP outperforms TTT-Linear, which in turn outperforms Mamba.
Similar to TTT-Linear, RWKV [55, 56], xLSTM [5], and Gated Linear Attention (GLA) [79] also have matrix hidden states, which are inherited from linear attention [41]. Modern RNNs such as GLA use chunk-wise parallelism to improve hardware efficiency, so tokens inside a chunk can be processed with matmuls instead of a cumsum. However, chunk-wise parallelism does not change the expressiveness of the model, since all temporal dependencies are still equivalent to a cumsum.
In contrast, mini-batch TTT allows more complex temporal dependencies across mini-batches. Each hidden state $W_{t}$ depends on previous $W$s within its mini-batch still through a cumsum, but depends on $W$s in previous mini-batches also through the gradient operator. As illustrated Figure LABEL:fig:batch-size, mini-batch TTT enables a trade-off between expressiveness and hardware efficiency, since a smaller batch size $b$ leads to better perplexity at the cost of higher latency. This trade-off is a unique and important feature of TTT. As shown in Table 1, the intermediate batch size $b=16$ significantly outperforms $b=T$ which is fully cumsum.
@@ 4.2.1 Test-Time Training
Conceptually, the biggest difference between our paper and prior work is that our reconstruction task is learned in an outer loop, instead of handcrafted with human priors. Follow-up work to TTT has explored applications such as robot manipulation [28] and locomotion [68], among others, that often require different designs for the self-supervised task.
@@ 4.2.2 Fast Weights
The general idea of fast weights is to update the parameters of a “fast” model on only the most relevant data, as opposed to the conventional practice of updating a “slow” model on all data [71]. This idea has existed since the 1980s [32]. The most relevant data can be the test instance itself, therefore TTT can be viewed as a special case of fast weights.
Prior work in fast weights usually avoids forming an explicit learning problem that optimizes some objective on data. For example, the update rule of Hebbian learning and Hopfield networks [35] simply adds $xx^{T}$ (or some variant thereof) [4] to the fast weights given each input $x$. In contrast, TTT embraces the idea of formulating an explicit learning problem, where the test instance is the target of generalization. Our update rule is also an explicit step of optimization.
The idea of fast weight programmers (FWPs) is to update the fast weights with a “slow” model [62]. Our inner-loop weights $W$ can be viewed as “fast” and outer-loop weights $\theta$ as “slow”. Therefore, networks containing TTT layers can be viewed as a special case of FWPs [43], similar to how TTT can be viewed as a special case of fast weights. The FWP with the Hebbian update rule above is equivalent to linear attention [60], therefore also to naive TTT-Linear with batch gradient descent.
The definition of FWPs is very broad. In fact, all networks with some gating mechanism, such as Transformers with SwiGLU blocks [63], can also be viewed as a special case of FWPs161616 Consider a simple gate $z=\sigma(\theta x)\odot(\theta^{\prime}x)$, where $x$ is the input, $z$ is the output, $\theta$ and $\theta^{\prime}$ are learnable weight matrices, $\odot$ is element-wise multiplication, and $\sigma$ is the sigmoid function. A well known interpretation is to view $W=\texttt{diag}(\theta^{\prime}x)$ as the fast weights controlled by slow weights $\theta^{\prime}$, then equivalently, $z=W\sigma(\theta x)$ is simply a two-layer MLP with fast weights [26]. . Recent work has been experimenting with FWPs for language modeling: Irie et al. [37] design “fast” networks with weights produced as output of a “slow” networks. Clark et al. [16] give a Transformer a final layer of fast weights, whose initialization is trained as slow weights. Our contribution relative to existing work on FWPs, again, is formulating an explicit learning problem for the update, which enables us to borrow tools from learning such as mini-batch and LN.
@@ 5 Discussion
The search space for effective instantiations inside this framework is huge, and our paper has only taken a baby step. Fortunately, if our perspective holds, then heuristics from regular training can transfer to test-time training, and search can be efficient. Next we outline some especially promising directions for future work.
@@
Why do we study TTT? First a more basic question: Why study AI? For some of us, AI is a playground to probe about the nature of human intelligence. Prior work often tries to model human learning with machine learning, where training is on a shuffled dataset with i.i.d. instances, and inference is on a separate test set. However, humans do not naturally learn with i.i.d. instances or have a train-test split. We believe that human learning has a more promising connection with TTT, our inner loop, whose data is a potentially very long sequence with strong temporal dependencies, and any piece of data can be used for both training and testing. This is why we study TTT.
@@ Appendix A Dual Form
The goal of this section is to derive the dual form for general MLPs of arbitrary depth, with nonlinear activations.
@@ A.1 Forward pass
$l=\frac{1}{2}\ell\left(W_{0}^{1},\dots,W_{0}^{K};\hat{X}\right)=\frac{1}{2}\big{\|}\hat{X}^{K+1}-Y\big{\|}_{F}^{2}=\sum_{t=1}^{b}l_{t},$
@@ A.3 Dual form
  • $\nabla_{Z^{k}}l=\sigma^{\prime}_{k}\left(\nabla_{\hat{X}^{k+1}}l\right)$
@@
  • $\bar{Z}^{k}=W^{k}\bar{X}^{k}-\nabla_{Z^{k}}l\cdot\texttt{mask}\left(\left(\hat{X}^{k}\right)^{T}\bar{X}^{k}\right)$
  • $\bar{X}^{k+1}=\sigma(\bar{Z}^{k})$

### v3

@@ Abstract
Self-attention performs well in long context but has quadratic complexity. Existing RNN layers have linear complexity, but their performance in long context is limited by the expressive power of their hidden states. We present a practical framework for instantiating sequence modeling layers with linear complexity and expressive hidden states. The key idea is to make the hidden state a machine learning model itself, and the update rule a step of self-supervised learning. Since the hidden state is updated by training even on test sequences, our layers are called Test-Time Training (TTT) layers. We consider two instantiations: TTT-Linear and TTT-MLP, whose hidden state is a linear model and a two-layer MLP respectively. We evaluate our instantiations at the scale of 125M to 1.3B parameters, comparing with a strong Transformer and Mamba, a modern RNN. Similar to Transformer, TTT-Linear and TTT-MLP can keep reducing perplexity by conditioning on more tokens, while Mamba cannot after 16k context. TTT-MLP still faces challenges in memory I/O, but shows larger potential in long context, pointing to a promising direction for future research.
@@
Figure 2: Comparing to Mamba, TTT-Linear and TTT-MLP have similar perplexity in 8k context (left) and better use of long context (right). Evaluations follow Kaplan et al. [36]. Left: Scaling trends on the Pile with 8k context, zoomed in between 350M and 1.3B parameters. Right: Similar to Transformer, TTT-Linear and TTT-MLP can keep reducing perplexity by conditioning on more tokens, while Mamba cannot after 16k context. All methods have matched training FLOPs as Mamba 1.4B.
@@ 1 Introduction
The difficulty with long context is inherent to the very nature of RNN layers: Unlike self-attention, RNN layers have to compress context into a hidden state of fixed size. As a compression heuristic, the update rule needs to discover the underlying structures and relationships among thousands or potentially millions of tokens. This need is inherently challenging. In this paper, we begin with the observation that self-supervised learning can compress a massive training set into the weights of a model such as an LLM, which often exhibits deep understanding about the semantic connections among its training data – exactly what we need from a compression heuristic.
@@
Motivated by this observation, we make the hidden state a machine learning model itself, and the update rule a step of self-supervised learning. Since the hidden state is updated by training even on test sequences, these RNN layers are called Test-Time Training (TTT) layers. We introduce two simple instantiations: TTT-Linear and TTT-MLP, where the hidden state is a linear model and a two-layer MLP, respectively. TTT layers can be integrated into any network architecture and optimized end-to-end, similar to RNNs layers and self-attention.
@@
We apply two techniques to make TTT layers more efficient on modern GPUs and TPUs. First, similar to the standard practice of taking gradient steps on mini-batches of sequences during regular training for better parallelism, we use mini-batches of tokens during TTT. Second, we develop a dual form for operations inside each TTT mini-batch. The dual form is equivalent in output to the naive implementation, but trains more than 5×5$\times$5 × faster on our TPUs.
Contributions and limitations.
The idea of using linear models as hidden states has already been well studied in DeltaNet [55, 75]. Since our first version was released, RNN layers with matrix (linear) hidden states have also been further advanced in Mamba 2 [17] and Gated DeltaNet [74]. Compared to this line of work, our contribution is a practical framework that can instantiate arbitrary neural networks as hidden states. However, such instantiations can still require substantial wall-clock time, even after applying our improvements in efficiency. It remains to be seen whether our framework can produce instantiations that either overcome this limitation or offer benefits outweighing it.
@@ 3.2 Long context: Books
To evaluate capabilities in long context, we experiment with context lengths ranging from 1k to 32k in 2×2$\times$2 × increments, using a popular subset of the Pile called Books3. The training recipe here is the same as that for Pile.121212 Following the Mamba paper, we always use 0.5M tokens per training batch regardless of the context length. In other words, for context length $T$ we have $0.5\text{M}\,/\,T$ sequences per batch (assume divisible). From the subset of results in Figure 11, we make a few observations:
@@
Experiments in Figure 2 (right).
Compared to TTT-Linear, TTT-MLP with matched FLOPs performs worse at short context but better at long context. This observation matches our expectation that the MLP as hidden state is more expressive than the linear model: The larger capacity of a more expressive hidden state is well-utilized in long context (therefore an advantage), but redundant in short context (therefore a disadvantage in our setting with matched FLOPs). The Transformer in this figure is TF finetune, which is the stronger baseline in 32k context. Details of the experiments in Figure 2 are included in Appendix C.
@@ 3.3 Wall-clock time
Figure 12: Latency on an NVIDIA A100 GPU with 80G HBM and PCIe connections.
@@
Due to resource constraints, our experiments are written in JAX and run on TPUs. On a v5e-256 TPU pod, the Transformer baseline takes 0.30s per iteration of training at context 2k, while TTT-Linear takes 0.27s per iteration, already 10% faster without any systems optimization. However, Mamba (implemented in PyTorch, Triton, and CUDA) can only run on GPUs, so for fair comparison, we also rewrite our method into GPU kernels. We only write inference kernels for this work because the training kernel would require substantial effort and cannot be used on our TPUs.
Figure 12 shows the latency of our inference kernel for forward (prefill) and generate (decode). All models are 1.3B (1.4B for Mamba). As expected, time per token grows linearly for Transformer as the context length increases, but stays roughly constant for the other methods.131313 We observe that the forward latency of the network increases slightly for TTT-Linear, TTT-MLP, and Mamba, even though the latency of each sequence modeling layer alone stays constant. Consider the operation $\theta X$, where $\theta$ is $d\times d$ and $X$ is $d\times T$. Its latency (normalized over $T$) is expected to be constant, but in practice grows slightly with $T$. One possible cause of this phenomenon is the GPU throttling after $T$ gets very large [28]. Note that our Transformer baseline is significantly faster that in the Mamba paper, because we use vLLM [42], a state-of-the-art serving system, instead of the HuggingFace Transformer [72].
@@ 4.1.1 Test-Time Training
Conceptually, the biggest difference between our paper and prior work is that our reconstruction task is learned in an outer loop, instead of handcrafted with human priors. Follow-up work has explored applications such as robot manipulation [26] and locomotion [64], among others, that often require different designs for the self-supervised task. In a preliminary manuscript [63], we explore the idea of learning to (learn at test time) from the perspective of TTT with reconstruction, with experiments in object recognition.
@@ 4.1.2 Fast Weights
The general idea of fast weights is to update the parameters of a “fast” model on only the most relevant data, as opposed to the conventional practice of updating a “slow” model on all data [67]. This idea has existed since the 1980s [30]. The most relevant data can be the test instance itself, therefore TTT can be viewed as a special case of fast weights. Compared to fast weights, TTT embraces the idea of formulating an explicit learning problem, where the test instance is the target of generalization. Our update rule is also an explicit step of optimization.
The idea of fast weight programmers (FWPs) is to update the fast weights with a “slow” model [58]. As a modern example for language modeling, Clark et al. [15] give a Transformer a final layer of fast weights, whose initialization is trained as slow weights. Our inner-loop weights $W$ can be viewed as “fast” and outer-loop weights $\theta$ as “slow”. Therefore, networks containing TTT layers can be viewed as a special case of FWPs [39], similar to how TTT can be viewed as a special case of fast weights.
Modern RNN layers such as linear attention [37, 56] and DeltaNet [55, 75] are inspired by the idea of FWPs. Given their relevance to our work, we discuss these modern RNN layers in detail in the next subsection.
4.2 Modern RNN layers
Our baseline, Mamba [25], is only one of the many recent RNN layers that inherit the linear (matrix) hidden states of linear attention [37, 56]. Some more recent examples are RWKV [51, 52], xLSTM [4], and Gated Linear Attention (GLA) [74]. The most relevant work is DeltaNet [55], which is equivalent to TTT-Linear with inner-loop mini-batch size 1, without the Layer Norm and residual connection. In [75], Yang et al. further improve the performance of DeltaNet and enable parallelized updates across tokens (in our terms, across inner loop mini-batches). Since our first version was released, RNN layers with matrix (linear) hidden states have also been further advanced in Mamba 2 [17] and Gated DeltaNet [74].
Compared to this line of work, our contribution is a practical framework that can instantiate arbitrary neural networks as hidden states. However, such instantiations can still require substantial wall-clock time, even after applying our improvements in efficiency. For example, TTT-MLP is effective in terms of FLOPs, as shown in Figure 2. But the additional complexity of the MLP structure increases wall-clock time much more relative to FLOPs, as shown in Figure 12. It remains to be seen whether our framework can produce instantiations that either overcome this limitation or offer benefits outweighing it.
@@ 5 Discussion
Future work.
The search space for effective instantiations inside this framework is huge, and our paper has only taken a baby step. Fortunately, if our perspective holds, then heuristics from regular training can transfer to test-time training, and search can be efficient. Next we outline some especially promising directions for future work:
@@
Why do we study TTT?
First a more basic question: Why study AI? For some of us, AI is a playground to probe about the nature of human intelligence. Prior work often tries to model human learning with machine learning, where training is on a shuffled dataset with i.i.d. instances, and inference is on a separate test set. However, humans do not naturally learn with i.i.d. instances or have a train-test split. We believe that human learning has a more promising connection with TTT, our inner loop, whose data is a potentially very long sequence with strong temporal dependencies, and any piece of data can be used for both training and testing. This is why we study TTT.
@@ Appendix A Dual Form
Here we derive the dual form for general MLPs of arbitrary depth, with nonlinear activations.
@@ A.1 Forward pass
$l=\frac{1}{2}\ell\left(W_{0}^{1},\dots,W_{0}^{K};X\right)=\frac{1}{2}\big{\|}\hat{X}^{K+1}-Y\big{\|}_{F}^{2}=\sum_{t=1}^{b}l_{t},$
@@ A.3 Dual form
  • $\nabla_{Z^{k}}l=\sigma^{\prime}_{k}\left(Z^{k}\right)\odot\nabla_{\hat{X}^{k+1}}l$
@@
where $\sigma^{\prime}$ is applied element-wise, and $\odot$ is element-wise multiplication.
@@
  • $\bar{Z}^{k}=W_{0}^{k}\bar{X}^{k}-\nabla_{Z^{k}}l\cdot\texttt{mask}\left(\left(\hat{X}^{k}\right)^{T}\bar{X}^{k}\right)$
  • $\bar{X}^{k+1}=\sigma\left(\bar{Z}^{k}\right)$